A very thing charged rod is placed along x axis with its midpoint at the origin.The charge per unit length of rod is Lemda c/m,the electric field vector at y=-a on u axis is
1
Expert's answer
2020-10-21T10:16:14-0400

The Gauss's theorem is given by a equation:

"div(E) = -4\pi\lambda"

"\int _V div (E) = \int_S E ds" - Gauss's theorem

"\int_V div (E) = -4\pi\lambda \int_L dl = -4\pi \lambda L"

"\int_S Eds = 2\pi |y|LE(x,y)"

If we've integrated it and evaluated E(x,y) then :

"E(x,a) = -2\frac{\lambda}{|a|}"